OpenStudy (anonymous):

2 more than a number equal to 30

OpenStudy (mathstudent55):

You are looking for an number. Let's call the unknown number x. 2 more than a number means 2 more than x. That means 2 added to x, or x + 2. We are told that that sum is equal to 30. The equation then is: x + 2 = 30 Now solve for x. You want x alone, so subtract 2 from both sides to get: x + 2 - 2 = 30 - 2 x = 28 The number is 28.